OpenStudy (anonymous):

Estimation.. the square root of 2?

OpenStudy (anonymous):

does. not. compute.

OpenStudy (anonymous):

by what method?

myininaya (myininaya):

\[\sqrt{1}<\sqrt{2}<\sqrt{4}\]

OpenStudy (anonymous):

i bet it is not newton's method that i for sure

myininaya (myininaya):

2 is closer to 1 \[1<\sqrt{2}<2\] so sqrt{2} is closer to 1 than it is 2

OpenStudy (anonymous):

two is closer to one? than what??

myininaya (myininaya):

so \[\sqrt{2} \approx 1.4\]

myininaya (myininaya):

no i said sqrt{2} is closer to 1 than it is 2

myininaya (myininaya):

and he knows no newton method or calculus ways only algebra satellite